AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 566     Entry time: Fri May 31 19:12:01 2024
Author: CB, EM 
Subject: Run 13 - test with beam 

20:00

The beam is set at the correct energy, operators cross-checked it again. What we see  at 750 is due to the molecular beam.  The operotors are trying to optimize and get rid of this in the next 30min.

Temps and system checks OK

Leakage current OK

Merger and tape server OK

Potentiometer at 28.06 mm and 29.23 mm

20:30

Keep going with the beam tests. A new peak appears around 950 keV( See attached). Still try to understand what it is. The operators keep going

Attachment 1: Screenshot_from_2024-05-31_20-13-46.png  602 kB  Uploaded Fri May 31 20:13:51 2024  | Hide | Hide all
Screenshot_from_2024-05-31_20-13-46.png
Attachment 2: Screenshot_from_2024-05-31_20-27-58.png  475 kB  Uploaded Fri May 31 20:28:06 2024  | Hide | Hide all
Screenshot_from_2024-05-31_20-27-58.png
Attachment 3: Screenshot_from_2024-05-31_20-32-59.png  589 kB  Uploaded Fri May 31 20:34:21 2024  | Hide | Hide all
Screenshot_from_2024-05-31_20-32-59.png
ELOG V3.1.4-unknown